Solve y^1=(2x+2y+3)/(2x+2y-1)

1
Expert's answer
2022-03-01T05:25:53-0500
"y'=\dfrac{2x+2y+3}{2x+2y-1}"

Let "x+y=t." Then

"t'_x=1+y'"

"t'-1=\dfrac{2t+3}{2t-1}"

"t'=\dfrac{4t+2}{2t-1}"

"\dfrac{2t-1}{4t+2}dt=dx"

Integrate


"\int \dfrac{2t-1}{4t+2}dt=\int dx"


"\int \dfrac{2t-1}{4t+2}dt=\dfrac{1}{2}\int \dfrac{4t+2}{4t+2}dt-\int \dfrac{1}{2t+1}dt"

"=\dfrac{1}{2}t-\dfrac{1}{2}\ln|2t+1|+C_1"

"\dfrac{1}{2}t-\dfrac{1}{2}\ln|2t+1|=x+\dfrac{1}{2}C"


"x+y-\ln|2x+2y+1|=2x+C"

"y-x-\ln|2x+2y+1|=C"